A point on the terminal side of an angle thetaθ in standard position is left parenthesis 2 comma negative 4 right parenthesis(2,− 4). Find the exact value of each of the six trigonometric functions of thetaθ.
Answer by stanbon(75887)

A point on the terminal side of an angle thetaθ in standard position is left parenthesis 2 comma negative 4 right parenthesis(2,− 4). Find the exact value of each of the six trigonometric functions of thetaθ.
-----------
tan(t) = -4/2 = -2/1
---
Since tan = y/x, y = -2 and x = 1
---
Then r = sqrt(2^2+1^2) = sqrt(5)
-----
sin = y/r = -2/sqrt(5)
cos = x/r = 1/sqrt(5)
tan = y/x = -2
--
csc = r/y = -sqrt(5)/2
sec = r/x = sqrt(5)
cot = x/y = -1/2
---------
Cheers,
Stan H.
------------